OpenStudy (anonymous):

The quadratic equation px^2+qx-36=0 has roots -2 and 6.Find a)the values of p and q.

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

so since -2, 6 are roots then (x+2)(x-6)=px^2+qx-36=0

OpenStudy (ikram002p):

expand (x+2)(x-6) can u do that ?

OpenStudy (anonymous):



OpenStudy (ikram002p):

oh wait sorry

OpenStudy (anonymous):

okay

OpenStudy (ikram002p):

the question has a mistake , it should be -12 not -36 but its ok it can be solved in anotherway

OpenStudy (anonymous):

k

OpenStudy (ikram002p):

so , -2 , and 6 are roots means f(-2)=f(-6)=0

OpenStudy (anonymous):

yes

OpenStudy (ikram002p):

f(x)=px^2+qx-36=0 f(-2)=4p-2q-36=0 .................... equation 1 f(6)=36 p +6q-36=0...................equation 2

OpenStudy (ikram002p):

could u solve the two linear equations now ?

OpenStudy (anonymous):

i will try :)

OpenStudy (ikram002p):

good , try and tell me what u got these are the equations 4p-2q=36 36p+6q=36

OpenStudy (anonymous):